In general, a statement involving n variables x1 , x 2 ,…, x n can be denoted by P( x1 , x 2 ,…, x n ) and is the value of the predicate P at the n-tuple (x1 , x2 ,…, x n ) . (a) Suppose Q(x , y) denotes the statement “ x=y+3 .” What are the truth values of the propositions Q(1,2) and Q(3,0)? (b) Let R(x , y ,


1
Expert's answer
2022-02-11T10:15:08-0500

(a) Suppose "Q(x , y)" denotes the statement “"x=y+3"”. Let us find the truth values of the propositions "Q(1,2)" and "Q(3,0)." Since "1\ne 5=2+3," we conclude that the proposition "Q(1,2)"is false. Taking into account that "3=0+3," we conclude that the proposition "Q(3,0)" is true.